OpenStudy (anonymous):

Aaron borrowed $200 to finance a VCR at a rate of 8.25% for 1.5 years. How much did he repay altogether?

OpenStudy (anonymous):

If you use: A=Ao(1+r)^t A=final amount Ao = initial amount (200) r= % (0.0825) t=time (1.5) If you plug those numbers in depending on how often they run the interest you should get your answer, unless it is compound interest which is a different equation